OpenStudy (across):

(Vector Calculus) A student asked me the following: \[\frac{d}{dt}\left[\vec{R}\cdot\vec{R}\right]=\frac{d\vec{R}}{dt}\cdot\vec{R}+\vec{R}\cdot\frac{d\vec{R}}{dt}=2\vec{R}\cdot\frac{d\vec{R}}{dt}=2\vec{R}\cdot\vec{V}.\]However (from dot-product properties), \[\frac{d}{dt}\left[\vec{R}\cdot\vec{R}\right]=\frac{d}{dt}\left[\left|\vec{R}\right|^2\right]=2\left|\vec{R}\right|\cdot\frac{d\vec{R}}{dt}=2\left|\vec{R}\right|\cdot\vec{V}.\]What's the catch here?

OpenStudy (amistre64):

its prolly in the 2(R.R')

OpenStudy (amistre64):

or not :)

OpenStudy (jamesj):

Note that in the second method it is \[2 |R| \frac{d |R|}{dt}\] Note also this isn't a dot product, just regular multiplication of real numbers or real-valued functions.

OpenStudy (across):

I was thinking perhaps I'm performing the chain rule incorrectly, as there's an absolute value, if I may, given these are vectors, in the expression.

OpenStudy (anonymous):

On the second line you can't have a scalar (|R|) dotted with a vector (V).

OpenStudy (across):

I know.

OpenStudy (across):

Apparently, \[2\vec{R}\cdot\vec{V}=2\left | \vec{R} \right |\frac{d\left | \vec{R} \right |}{dt}?\]

OpenStudy (jamesj):

Let's do an example: uniform circular motion, R = (cos t, sin t) Then 2R.R' = 2(cos t, sin t).(-sin t, cos t) = 0 and 2|R| d|R|/dt = 2 times 1 times 0 = 0, as |R| = 1 for all t So the two formulae are consistent.

OpenStudy (jamesj):

Yes.

OpenStudy (jamesj):

It makes sense physical this way. Write R is polar coordinates (or spherical polars if you're in R^3) Then the dot product tells you how much the velocity is in the direction of the radial displacement, as the spherical terms dot product to zero

OpenStudy (phi):

I'm thinking R and V are orthogonal?

OpenStudy (jamesj):

and then the other expression is exactly the radial displacement times the rate of change of the radial displacement. So the two interpretations are the same thing.

OpenStudy (jamesj):

In the example above, of course there is no change in radial displacement, so both terms are zero.

OpenStudy (phi):

In the second equation, once you do R dot R don't you get a scalar, and the derivative is zero.

OpenStudy (jamesj):

no, because it's a function of t